Calculate the energy associated with the magnetic field of a 188-turn solenoid in which a current of 1.48 A produces a magnetic flux of 3.81 10-4 T
Answer #1

Energy = (1/2) volume * B^2 / u

Flux = B * Area

B = u N i / length for a solenoid

and volume = length * area

putting this all together... eliminating volume, area, length, B by substitution

Energy = (1/2) * length * area (uNi /length )^2 / u

and flux = uNi * area / length

or

Energy = (1/2) u N^2 i^2 area / length = (1/2) flux * N * i =

= (1/2) * 0.000381 * 188 * 1.48 =

= 0.0530 Joules